9. Suppose the firms production function is given by f(K,L) min (K,L (a) For what values of a will the firm exhibit decrea
Answer #1

9. a. For the production function to exhibit decreasing returns to scale, increasing inputs by \small \beta times increases the output by less than \small \beta times. lal, ak) = min (3 ),(8L) = 3 minik, 2→act
For the production function to exhibit increasing returns to scale, increasing inputs by \small \beta times increases the output by greater than \small \beta times. lal, ak) = min (3 ),(8L) = 3 mink, 29→a>
For the production function to exhibit constant returns to scale, increasing inputs by \small \beta times increases the output by \small \beta times. lal, ak) = min (3 ),(8L) = 3 minik, 2)→ a=1

b. The producer treats the two inputs as perfect complements and always employs them in the fixed ratio:
Lº=9 L =q , K = 9K= 1/
The firm's cost function is:
Cost = wqa + role = (w+ng/

c. The new production function is:
(L,K) = min100,21
b=7+b= 17
Since the firm wants to produce less than 100 units of output, it will employ less than 10,000 units of labor.